The sum of two numbers is 18. The difference between four times the smaller number and seven is
equal to the sum of two times the larger number and five. Find the two numbers.

Answer by checkley77(12569)

x+y=18 or x=18-y
4x-7=2y+5
4(18-y)-7=2y+5
72-4y-7=2y+5
-4y-2y=5-72+7
-6y=-60
y=-60/-6
y=10 ans.
x=18-10
x=8 ans.
Proof:
4*8-7=2*10+5
32-7=20+5
25=25